# Entropy Scaling for Diffusion Coefficients in Mixtures

### v1

@@ Abstract
Entropy scaling is a powerful technique that has been used successfully for predicting transport properties of pure components over a wide range of states. However, modeling mixture diffusion coefficients by entropy scaling is an unresolved task. We have tackled this issue and present an entropy scaling framework for predicting mixture self-diffusion as well as mutual diffusion in a thermodynamically consistent way. The predictions of the mixture diffusion coefficients are made based on information on the self-diffusion coefficients of the pure components and the infinite dilution diffusion coefficients. This is accomplished using information on the entropy of the mixture, which is taken here from molecular-based equations of state. The framework is flexible and can be applied together with any equation of state. Examples for the application of the new entropy scaling framework for the prediction of diffusion coefficients in mixtures illustrate its impressive performance. It enables predictions a wide range of temperatures and pressures including gaseous, liquid, supercritical, and metastable states - also for strongly non-ideal mixtures.
I Introduction
Diffusion in mixtures is important in many natural and technical processes. And, numbers for diffusion coefficients are needed in many applications, e.g. for the design of fluid separation processes and reactors or for modeling combustion processes. However, experimental data on diffusion coefficients are notoriously scarce, so that reliable models for their prediction are needed. We introduce a new model type that is based on entropy scaling and enables predictions that were previously infeasible.
In diffusion, two phenomena are distinguished: Self-diffusion (a.k.a. tracer diffusion) and mutual (transport) diffusion. Self-diffusion and the corresponding self-diffusion coefficient $D_{i}$ of a component $i$ describes the Brownian movement of individual particles and is defined for pure components and mixtures. In contrast, mutual diffusion is only defined in mixtures and describes the motion of particle collectives of the different components that results in macroscopic mass transfer. Obviously, self-diffusion and mutual diffusion are closely related, but there exist no generally applicable equations that relate self-diffusion coefficients and mutual diffusion coefficients. The different diffusion coefficients are schematically depicted in Fig. 1 for a binary mixture.
Figure 1: Schematic representation of the diffusion coefficients in a binary mixture as a function of the mole fraction. Lines indicate the self-diffusion coefficient of component $i$ $D_{i}$ (blue) and component $j$ $D_{j}$ (green), the Maxwell-Stefan diffusion coefficient Ð$\textit{\DH}_{ij}$ (dark grey), and the Fickian diffusion coefficient $D_{ij}$ (light grey). The limiting cases $D_{i}^{\rm pure}$, $D_{j}^{\rm pure}$ (both pure cases) and $D_{i}^{\infty}$, $D_{j}^{\infty}$ (both pseudo-pure cases) are also depicted.
@@ I Introduction
$D_{ij}=\textit{\DH}_{ij}\Gamma_{ij},$ (1)
where $\Gamma_{ij}$ is the thermodynamic factor, which is defined by
$\Gamma_{ij}=\frac{x_{i}x_{j}}{RT}\frac{\partial^{2}G}{\partial x_{i}^{2}},$ (2)
@@
Evidently, the two coefficients $D_{ij}$ and Ð$\textit{\DH}_{ij}$ become equal if the thermodynamic factor is unity, which is the case in the infinite dilution limit (and for ideal mixtures). Furthermore, self-diffusion coefficient and the mutual diffusion coefficients are related in the infinite dilution limit (cf. Fig. 1). Thus, the following relations apply for the infinite dilution limit:
@@
where $D_{i}^{\infty}$ is the diffusion coefficient of component $i$ infinitely diluted in component $j$, $D_{j}^{\infty}$ the diffusion coefficient of component $j$ infinitely diluted in component $i$. Modeling the different diffusion coefficients in a mixture in a consistent way is a challenging task. In this work, we propose a new methodology that provides such a framework.
@@
Physical models for predicting mixture diffusion coefficients at gaseous states are known and established for a long time within kinetic gas theory [2] - see Supplementary Information for details. The prediction of mixture diffusion coefficients at states where significant intermolecular interactions prevail, on the other hand, is still an unresolved problem. Einstein has proposed a method for estimating infinite dilution diffusion coefficients in liquids [3], for which several modifications exist today [2, 4, 5]. For estimating the concentration dependence of mutual diffusion coefficients in mixtures, several empirical models have been proposed, e.g. the Vignes model and the generalized Darken model [6, 7, 8]. However, these models often fail for strongly non-ideal mixtures (see Supplementary Information).
@@
In recent years, entropy scaling has received significant attention for modeling transport properties [9, 10, 11]. It is based on the discovery of Rosenfeld [12, 13], that dynamic properties (i.e. viscosity, thermal conductivity, and self-diffusion coefficient) of pure components, when properly scaled by the density and temperature, are a monovariate function of the configurational entropy $s_{\rm conf}$. This scaling behavior is physically based and related to isomorph theory [14, 11] (however, it is not yet fully understood). For describing the entropy with respect to the desired state point (e.g. given by $T,p$), usually an equation of state (EOS) is used. Entropy scaling can be favorably combined with molecular-based EOS[15], which enables predictions beyond the available data [16, 17]. Entropy scaling is well established for predicting the viscosity and thermal conductivity of mixtures [18, 19]. The corresponding models are based on combination and mixing rules and often enable a reliable prediction of mixture viscosities and thermal conductivities. However, entropy scaling for mixture diffusion coefficients – as depicted in Fig. 1 – has not yet been developed. So far, only the pure component limiting cases for the self-diffusion coefficients $D_{1}^{\rm pure}$ and $D_{2}^{\rm pure}$ (cf. Fig. 1) can be described by entropy scaling models from the literature[20, 21, 22, 23, 16]. Molecular dynamics simulation results indicate that also mixture diffusion coefficients might be subject to a scaling with respect to the entropy [24, 10, 25]. However, no modeling framework for describing mixture diffusion coefficients has been developed yet.
@@
In this work, we propose an entropy scaling model for predicting mixture diffusion coefficients, namely the self-diffusion coefficients as well as the (Fickian and MS) mutual diffusion coefficients, without any adjustable mixture parameters. The physical framework we have developed for predicting diffusion coefficients in mixtures is unique and can be characterized as follows: (1) It can be applied in the entire fluid region, i.e. it covers gases, liquids, and supercritical fluids, phase equilibria, and even metastable states. (2) It describes both self-diffusion and mutual diffusion in a consistent way. (3) It comprises the dependence of the diffusion coefficients on temperature, pressure and composition in the entire fluid region. The applicability of the model is demonstrated for binary mixtures in this work.
@@
The approach developed in this work is based on three central ideas and concepts: (1) Infinite dilution diffusion coefficients are treated as pseudo-pure components that exhibit a monovariate scaling behavior, which can be treated by classical entropy scaling [16]. (2) Therefore, the mixture diffusion coefficient limiting cases, i.e. the pure component and pseudo-pure component self-diffusion coefficients (cf. Fig. 1), are modeled as functions of the entropy. This requires at least one reference data point for each limiting case. (3) Based on the information of the limiting cases, the concentration dependence of the diffusion coefficients $D_{i}$, $D_{j}$, Ð$\textit{\DH}_{ij}$, and $D_{ij}$ are predicted using combination and mixing rules. We demonstrate that these predictions can succeed without any adjustable mixture parameter. The performance of the new approach is demonstrated using model fluids as well as real substance systems. In this work, we use molecular-based EOS models. Yet, also other EOS types such as multiparameter [26] or cubic EOS [27] could be used.
II Results
@@
First, we demonstrate the applicability of a monovariate scaling to the infinite dilution self-diffusion coefficients by treating $D_{i}^{\infty}$ as a pseudo-pure component. This enables the prediction of $D_{i}^{\infty}$ to practically all fluid states based on very little data. Then, we present results for the prediction of the different mixture diffusion coefficients for selected examples (more given in the Supplementary Information).
II.1 Infinite Dilution Diffusion Coefficients
@@
Fig. 2a demonstrates the monovariate scaling behavior (cf. Eqs. (5) - (7)) for infinite dilution diffusion coefficients $\widehat{D}_{2}^{\infty,\circ}$ for a binary Lennard-Jones mixture.
@@
Figure 2: Entropy scaling of diffusion coefficients in a Lennard-Jones mixture with $\sigma_{2}=\sigma_{1}$, $\varepsilon_{2}=0.9\,\varepsilon_{1}$, and $\varepsilon_{12}=1.2\,\sqrt{\varepsilon_{1}\varepsilon_{2}}$. a) Scaled infinite dilution diffusion coefficient of component 2 $\widehat{D}^{\infty,\circ}_{2}$ as a function the reduced configurational entropy $s_{\rm conf}/R$. The line indicates the entropy scaling model. Symbols are MD simulation data from this work. The inset shows the simulation state points in the temperature-density phase diagram of component 1. Therein, solid lines indicate the phase envelopes from Refs. 28, 29. b) Self-diffusion coefficient of component 2 $D_{2}$ (top) and infinite dilution diffusion coefficient of component 2 $D^{\infty}_{2}$ (bottom) as function of the pressure. Lines are the entropy scaling model and symbols are simulation results from Ref. 16 ($D_{2}^{\rm pure}$) and from this work ($D^{\infty}_{2}$). The black solid lines indicates the vapor-liquid equilibrium and the black dashed line the corresponding spinodal. Colors indicate the temperature.
@@
This is impressive considering the fact that a large range of states was studied, cf. Fig. 2a-inset. The quality of the scaling of $\widehat{D}_{2}^{\infty,\circ}$ (i.e. how well the data can be described by a monovariate function) is essentially the same as that found for the pure component diffusion coefficient $D_{2}$. This supports the picture introduced in this work that $D_{2}^{\infty}$ can be considered as a pseudo-pure component property. This picture is physically meaningful considering the fact that in both cases, the mobility of a single particle in a homogeneous environment is described. The Chapman-Enskog diffusion coefficient agrees well with the low-density simulation results – as indicated by the convergence of $\widehat{D}^{\infty,\circ}_{2}\rightarrow 1$ for $s_{\rm conf}/R\rightarrow 0$. In the Supplementary Information, the results for a second binary Lennard-Jones system and a real substance system are presented that support the findings discussed here. For the two Lennard-Jones model systems, the "global" parameters from Ref. 16 were sufficient for describing $\widehat{D}_{2}^{\infty,\circ}$. This finding was not expected and demonstrates the broad applicability of the universal parameters $g_{1}$ and $g_{2}$ in the correlation, that were determined in Ref. 16.
@@
From the models for the pure component self-diffusion coefficient $\widehat{D}_{2}^{\rm pure}(\tilde{s}_{\rm conf})$ [16] and the infinite dilution diffusion coefficient $\widehat{D}_{2}^{\infty}(\tilde{s}_{\rm conf})$, both $D_{2}^{\rm pure}$ and $D_{2}^{\infty}$ can be predicted in a wide range of states. For the Lennard-Jones mixture, the configurational entropy was taken from the Kolafa-Nezbeda EOS[30]. Fig. 2b shows the results of the entropy scaling models for $D_{2}$ and $D_{2}^{\infty}$ in comparison to the MD results. The agreement between both methods is excellent, which is a result of the good performance of the EOS as well as of the fact that both properties show a monovariate relation with respect to the entropy.
@@
II.2 Diffusion Coefficients in Mixtures
@@
The model proposed in this work (see methods section) is able to predict all diffusion coefficients in a mixture, which is demonstrated here for binary systems. Fig. 3 shows the model predictions for all four diffusion coefficients ($D_{1}$, $D_{2}$, Ð$\textit{\DH}_{12}$, $D_{12}$) of the Lennard-Jones mixture for the entire composition range and different temperatures. For comparison, simulation data from Ref. [31] are used. The predictions from the entropy scaling framework and the computer experiment results agree well for all four diffusion coefficients (see Fig. 3).
@@
For both self-diffusion coefficients, for which the simulation results have a smaller statistical uncertainty than for the mutual diffusion coefficients, the entropy scaling model describes most simulation data within their uncertainty, despite the strong non-ideality that leads to extrema at about $x_{2}\approx 0.4\,{\rm mol\,mol^{-1}}$. The simulation data for Ð$\textit{\DH}_{12}$ scatter more and have larger error bars than those for $D_{1}$ and $D_{2}$. The entropy scaling framework predicts the Ð$\textit{\DH}_{12}$ simulation data very well in the entire composition range. In particular, the strongly non-ideal behavior is captured accurately by the entropy scaling framework in contrast to established empirical models such as the Vignes or Darken model that fail here (see Supplementary Information). Also the temperature and pressure dependency of the diffusion coefficients are very well described by the model (cf. Fig. 3 and Supplementary Information, respectively). For the predictions of the Fickian diffusion coefficient, the trends are correctly predicted by the entropy scaling predictions, but deviations are observed, cf. Fig. 3d. The deviations between simulation results and entropy scaling are larger for the Fickian diffusion coefficient, which can be attributed to the inclusion of the thermodynamic factor adding additional complexity to the model predictions (and the MD sampling [25]). In contrast to the other three diffusion coefficients ($D_{1}$, $D_{2}$, Ð$\textit{\DH}_{12}$), the Fickian diffusion coefficient $D_{12}$ exhibits a maximum at about $x_{2}\approx 0.55\,{\rm mol\,mol^{-1}}$, which is in line with values of $\Gamma_{12}\gg 1$ obtained in this system. Due to the coupling of entropy scaling and EOS, the model is inherently consistent and can be applied over a wide range of states. As an example, the results for a Lennard-Jones mixture with $\xi_{12}<1$ that exhibits a liquid-liquid miscibility gap are shown in the Supplementary Information.
@@
Figure 3: Diffusion coefficients in a binary Lennard-Jones mixture with $\sigma_{2}=\sigma_{1}$, $\varepsilon_{2}=0.9\,\varepsilon_{1}$, and $\varepsilon_{12}=1.2\,\sqrt{\varepsilon_{1}\varepsilon_{2}}$ as a function of the mole fraction $x_{2}$ at $p=0.13\,\sigma^{3}_{1}\varepsilon_{1}^{-1}$. a) Self-diffusion coefficient of component 1 $D_{1}$; b) Self-diffusion coefficient of component 2 $D_{2}$; c) Maxwell-Stefan diffusion coefficient Ð$\textit{\DH}_{12}$; d) Fickian diffusion coefficient $D_{12}$. Lines are predictions from the entropy scaling framework. Symbols are simulation results from Ref. 31. The entropy scaling framework was used in combination with the Kolafa-Nezbeda EOS[30]. The colors indicate the temperature $T\in\{0.79,0.855,0.92,0.985,1.05\}\,k_{\rm B}\varepsilon_{1}^{-1}$ (from blue to yellow).
@@
Fig. 4 shows the results for the diffusion coefficients in two real substance systems: A mixture of alkanes, $n$-hexane (C6H14) + $n$-dodecane (C12H26), and the mixture toluene (C7H8) + $n$-hexane (C6H14). For both systems, only little experimental data are available (which is common for practically all real mixtures), which is used for the validation of the predictions. The thermodynamic properties of the systems were modeled by the PC-SAFT EOS [32] and the component-specific models for the pure substances were taken from the literature (see Supplementary Information for details).
@@
Figure 4: Diffusion coefficients of real substance mixtures predicted by entropy scaling. Symbols are experimental data and lines are predictions from the entropy scaling framework. a) Self-diffusion coefficients $D_{1}$ in the mixture $n$-hexane (1) + $n$-dodecane (2) as a function of the mole fraction of $n$-dodecane $x_{2}$ at $p=0.1\,{\rm MPa}$. Experimental data are from Ref. 33. b) Fickian diffusion coefficient $D_{\rm 12}$ in the mixture toluene (1) + $n$-hexane (2) as a function of the mole fraction of $n$-hexane $x_{2}$ at $p=0.1\,{\rm MPa}$. Experimental data are from Ref. 34.

### v2

# Entropy scaling for diffusion coefficients in fluid mixtures

@@ Abstract
Entropy scaling is a powerful technique that has been used for predicting transport properties of pure components over a wide range of states. However, modeling mixture diffusion coefficients by entropy scaling is an unresolved task. We tackle this issue and present an entropy scaling framework for predicting mixture self-diffusion coefficients as well as mutual diffusion coefficients in a thermodynamically consistent way. The predictions of the mixture diffusion coefficients are made based on information on the self-diffusion coefficients of the pure components and the infinite-dilution diffusion coefficients. This is accomplished using information on the entropy of the mixture, which is taken here from molecular-based equations of state. Examples for the application of the entropy scaling framework for the prediction of diffusion coefficients in mixtures illustrate its performance. It enables predictions over a wide range of temperatures and pressures including gaseous, liquid, supercritical, and metastable states – also for strongly non-ideal mixtures.
keywords:
thermodynamics, diffusion, entropy scaling
Introduction
Diffusion in mixtures is important in many natural and technical processes. Numerical values for diffusion coefficients are needed in many applications, e.g., for the design of fluid separation processes, reactors, and combustion processes. However, experimental data on diffusion coefficients are notoriously scarce, so that reliable models for their prediction are needed. We introduce a model type that is based on entropy scaling and enables predictions that were previously infeasible.
In diffusion, two phenomena are distinguished: self-diffusion (a.k.a. tracer diffusion) and mutual (transport) diffusion. Self-diffusion and the corresponding self-diffusion coefficient $D_{i}$ of a component $i$ describes the Brownian movement of individual particles and is defined for pure components and mixtures. In contrast, mutual diffusion is only defined in mixtures and describes the motion of particle collectives of the different components resulting in macroscopic mass transfer. Obviously, self-diffusion and mutual diffusion are closely related, but there exist no generally applicable relation that connects self-diffusion coefficients and mutual diffusion coefficients. The different diffusion coefficients are schematically depicted in Fig. 1 for a binary mixture.
Figure 1: Schematic representation of the different diffusion coefficients in a binary mixture as a function of the mole fraction. Lines indicate the self-diffusion coefficient of component $i$ $D_{i}$ (blue) and component $j$ $D_{j}$ (green), the Maxwell-Stefan diffusion coefficient Ð$\textit{\DH}_{ij}$ (dark grey), and the Fickian diffusion coefficient $D_{ij}$ (light grey). The limiting cases $D_{i}^{\rm pure}$, $D_{j}^{\rm pure}$ (both pure cases) and $D_{i}^{\infty}$, $D_{j}^{\infty}$ (both pseudo-pure cases) are also depicted.
@@ Introduction
$D_{ij}=\textit{\DH}_{ij}\varGamma_{ij},$ (1)
where $\varGamma_{ij}$ is the thermodynamic factor, which is defined by
$\varGamma_{ij}=\frac{x_{i}x_{j}}{{\rm R}T}\left(\frac{\partial^{2}G}{\partial x_{i}^{2}}\right)_{T,p,n_{j\neq i}},$ (2)
@@
The two coefficients $D_{ij}$ and Ð$\textit{\DH}_{ij}$ become equal if the thermodynamic factor is unity, which is the case in the infinite dilution limit (and for ideal mixtures). Furthermore, the self-diffusion coefficient and the mutual diffusion coefficients are related in the infinite dilution limit (cf. Fig. 1). Thus, the following relations apply for the infinite dilution limit:
@@
where $D_{i}^{\infty}$ is the diffusion coefficient of component $i$ infinitely diluted in component $j$, $D_{j}^{\infty}$ the diffusion coefficient of component $j$ infinitely diluted in component $i$. Modeling the different diffusion coefficients in a mixture in a consistent way is a challenging task. In this work, we propose a methodology that provides such a framework.
@@
Physical models for predicting mixture diffusion coefficients at gaseous states are known and established for a long time within kinetic gas theory [2]. The prediction of mixture diffusion coefficients at states where significant intermolecular interactions are present, on the other hand, is still an unresolved problem. Einstein has proposed a method for estimating infinite-dilution diffusion coefficients in liquids [3], for which several modifications exist today [2, 4, 5]. For estimating the concentration dependence of mutual diffusion coefficients in mixtures, several empirical models have been proposed, e.g., the Vignes model and the generalized Darken model [6, 7, 8]. However, these models often fail for strongly non-ideal mixtures (see Suppl. Note 8).
@@
In recent years, entropy scaling has received significant attention for modeling transport properties [9, 10, 11]. It is based on the discovery of Rosenfeld [12, 13], that dynamic properties (i.e. viscosity, thermal conductivity, and self-diffusion coefficient) of pure components, when properly scaled by the density and temperature, are a monovariate function of the configurational entropy (sometimes also referred to as ’residual entropy’ or ’excess entropy’). This scaling behavior is physically based and related to isomorph theory [14, 11]. After being re-discovered by Dyre in a seminal review [11], entropy scaling has become a popular approach in the last ten years. For modeling transport properties, the entropy scaling principle has been cast into many executable models using several different modified scaling approaches [9, 15, 16, 17, 18, 19]. For obtaining the entropy at a desired state point (e.g., given by $T,p$), usually an equation of state (EOS) is used. Entropy scaling can be favorably combined with molecular-based EOS[20], which enables predictions beyond the available data [17, 21]. Entropy scaling is well established for predicting the viscosity and thermal conductivity of mixtures, as for example demonstrated by Gross and co-workers [9, 22] and Bell and co-workers [23, 24] in recent years. The corresponding models are based on combination and mixing rules and often enable a reliable prediction of mixture viscosities and thermal conductivities [25, 23, 17]. However, entropy scaling for mixture diffusion coefficients – as depicted in Fig. 1 – has not yet been developed. So far, only the pure component limiting cases for the self-diffusion coefficients $D_{1}^{\rm pure}$ and $D_{2}^{\rm pure}$ (cf. Fig. 1) can be described by entropy scaling models from the literature[26, 18, 19, 16, 17]. There are few molecular simulation studies of self-diffusion coefficients in mixtures available in the literature [27, 28, 29, 30, 31, 32, 10, 33] in which also the entropy scaling behavior of the data was considered. They evaluate either simple model systems or special cases like metallic fluids. They show that self-diffusion coefficients of such fluids can follow a universal monovariate behavior, but the authors do not provide models for describing the self-diffusion coefficients. Specifically, a quasi-universal scaling law for mixture diffusion coefficients has been proposed by Bell and Dyre [10] – however, being limited to self-diffusion. Truskett and co-workers [29, 30, 31, 32] have studied the monovariate scaling behavior based on computer experiment data. Similarly, Fertig et al.[33] have demonstrated that elements of monovariate scaling are also present in model mixture diffusion coefficient data. However, in none of these studies a generally applicable modeling framework for consistently describing the different diffusion coefficients in mixtures has been developed, so that this is still an unresolved issue.
@@
In this work, we propose an entropy scaling model for predicting mixture diffusion coefficients, namely the self-diffusion coefficients as well as the (Fickian and MS) mutual diffusion coefficients, without any adjustable mixture parameters. The physical framework we have developed for predicting diffusion coefficients in mixtures can be characterized as follows: (1) It can be applied in the entire fluid region, i.e. it covers gases, liquids, and supercritical fluids, phase equilibria, and even metastable states. (2) It describes both self-diffusion and mutual diffusion in a consistent way. (3) It comprises the dependence of the diffusion coefficients on temperature, pressure, and composition in the entire fluid region. The applicability of the model is demonstrated for binary mixtures in this work.
@@
The approach developed in this work is based on three central ideas and concepts: (1) infinite-dilution diffusion coefficients are treated as pseudo-pure components that exhibit a monovariate scaling behavior, which can be treated by classical entropy scaling [17]. (2) Therefore, the mixture diffusion coefficient limiting cases, i.e. the pure component and pseudo-pure component self-diffusion coefficients (cf. Fig. 1), are modeled as functions of the entropy. This requires at least one reference data point for each limiting case. (3) Based on the information of the limiting cases, the concentration dependence of the diffusion coefficients $D_{i}$, $D_{j}$, Ð$\textit{\DH}_{ij}$, and $D_{ij}$ are predicted using combination and mixing rules. (4) It correctly captures the limits of the different diffusion coefficients to the pure component and infinite dilution. We demonstrate that these predictions can succeed without any adjustable mixture parameter. The performance of the approach is demonstrated using model fluids as well as real substance systems. In this work, we employ molecular-based EOS models. Yet, also other EOS types such as multiparameter [34] or cubic EOS [35] could be used.
Results
@@
First, we demonstrate the applicability of a monovariate scaling to the infinite dilution self-diffusion coefficients by treating $D_{i}^{\infty}$ as a pseudo-pure component property. This enables the prediction of $D_{i}^{\infty}$ to practically all fluid states based on very little data. The monovariate scaling of the infinite-dilution diffusion coefficients uncovered in this work provides the basis for the modeling of the different mixture diffusion coefficients shown in the second part of this section. The applicability of the modeling approach is demonstrated in the second part of this section, where model predictions for the self-diffusion as well as mutual diffusion coefficients are compared to reference data for model and real substance systems.
Infinite-Dilution Diffusion Coefficients
@@
Fig. 2a demonstrates the monovariate scaling behavior for infinite-dilution diffusion coefficients $\widehat{D}_{2}^{\infty,\circ}$ for a binary Lennard-Jones mixture.
@@
Figure 2: Entropy scaling of diffusion coefficients in a Lennard-Jones mixture with $\sigma_{2}=\sigma_{1}$, $\varepsilon_{2}=0.9\,\varepsilon_{1}$, and $\varepsilon_{12}=1.2\,\sqrt{\varepsilon_{1}\varepsilon_{2}}$. a) Scaled infinite-dilution diffusion coefficient of component 2 $\widehat{D}^{\infty,\circ}_{2}$ as a function the reduced configurational entropy $s_{\rm conf}/{\rm R}$. The line indicates the entropy scaling model. Symbols are MD simulation data from this work. The inset shows the simulation state points in the temperature-density phase diagram of component 1. Therein, solid lines indicate the phase envelopes from Refs. [36, 37]. b) Self-diffusion coefficient of component 2 $D_{2}^{\rm pure}$ (top) and infinite-dilution diffusion coefficient of component 2 $D^{\infty}_{2}$ (bottom) as function of the pressure. Lines are the entropy scaling model and symbols are simulation results from Ref. [17] ($D_{2}^{\rm pure}$) and from this work ($D^{\infty}_{2}$). The black solid line indicates the vapor-liquid equilibrium and the black dashed line the corresponding spinodal. Colors indicate the temperature. Source data are provided as a Source Data file.
@@
Results are shown for the scaled self-diffusion coefficient of component 2 at infinite dilution $\widehat{D}_{2}^{\infty,\circ}$; in our notation, the hat $\widehat{}$ and refer to specific parts of the scaling, see Methods. The data collapse to a monovariate curve, which is impressive considering the fact that a large range of states was studied, cf. Fig. 2a-inset. The quality of the scaling of $\widehat{D}_{2}^{\infty,\circ}$ (i.e. how well the data can be described by a monovariate function) is essentially the same as that found for the pure component diffusion coefficient $D_{2}$ (see Suppl. Note 5 for quantification). This supports the picture introduced in this work that $D_{2}^{\infty}$ can be considered as a pseudo-pure component property. This picture is physically meaningful considering the fact that in both cases, the mobility of a single particle in a homogeneous environment is described. The Chapman-Enskog diffusion coefficient agrees well with the low-density simulation results – as indicated by the convergence of $\widehat{D}^{\infty,\circ}_{2}\rightarrow 1$ for $s_{\rm conf}/{\rm R}\rightarrow 0$.
Fig. 3 shows the simulation results for the infinite-dilution diffusion coefficients in the systems acetone + isobutane and ethanol + chlorine.
Figure 3: Scaling behavior of the infinite-dilution diffusion coefficients in the system acetone (1) + isobutane (2) (left) and ethanol (1) + chlorine (2). a,c) Studied state points (symbols) in the temperature-density phase diagram of the solvents. The line indicates the vapor-liquid equilibrium and the star the critical point as calculated from the PC-SAFT EOS [38]. b,d) Scaled infinite-dilution diffusion coefficient of the solutes $\widehat{D}_{\rm 1}^{\infty,\circ}$ as a function of the reduced configurational entropy $\tilde{s}_{\rm conf}$. The symbols are the simulation results (color indicates the temperature) and the line the entropy scaling model (fitted to the simulation results). The error bars represent the uncertainties of the molecular simulation results. Source data are provided as a Source Data file.
The simulations comprise states in the gas, liquid, and supercritical regions. Details on the simulations are provided in Suppl. Note 2. For both systems, the infinite-dilution diffusion coefficient shows a monovariate behavior with respect to the configurational entropy confirming the results for the Lennard-Jones model system. Significant deviations from the monovariate behavior are observed for the real systems only for $\tilde{s}_{\rm conf}$, i.e. for low-density state points. This might be due to larger uncertainties of the simulations in this region.
In Suppl. Note 6), the results for two more binary Lennard-Jones systems and a third real substance system are presented that support the findings discussed here. For the Lennard-Jones model systems, the "global" parameters from Ref. [17] were sufficient for describing $\widehat{D}_{2}^{\infty,\circ}$. This finding was not expected and demonstrates the broad applicability of the universal parameters $g_{1}$ and $g_{2}$ in the correlation that were determined in Ref. [17].
@@
From the models for the pure component self-diffusion coefficient $\widehat{D}_{2}^{\rm pure}(\tilde{s}_{\rm conf})$ [17] and the infinite-dilution diffusion coefficient $\widehat{D}_{2}^{\infty}(\tilde{s}_{\rm conf})$, both $D_{2}^{\rm pure}$ and $D_{2}^{\infty}$ can be predicted in a wide range of states. For the Lennard-Jones systems, the configurational entropy was taken from the Kolafa-Nezbeda EOS[39], which is known to give an accurate the description of Lennard-Jones fluids [40, 41]. We used the mixture implementation for the Kolafa-Nezbeda EOS from Refs. [40, 17]. Fig. 2b shows the results of the entropy scaling models for $D_{2}$ and $D_{2}^{\infty}$ in comparison to the MD results. Both methods are in good agreement, which is a result of the good performance of the EOS as well as of the fact that both properties show a monovariate relation with respect to the configurational entropy.
@@
Diffusion Coefficients in Mixtures
@@
The model proposed in this work (see Methods) is able to predict all diffusion coefficients in binary mixtures. Fig. 4 shows the model predictions for all four diffusion coefficients ($D_{1}$, $D_{2}$, Ð$\textit{\DH}_{12}$, $D_{12}$) of three Lennard-Jones systems for the entire composition range and different temperatures. For comparison, simulation data from Ref. [42] are used. The predictions from the entropy scaling framework and the computer experiment results agree well for all investigated systems and all four diffusion coefficients (see Fig. 4).
@@
For both self-diffusion coefficients, for which the simulation results have a smaller statistical uncertainty than for the mutual diffusion coefficients, the entropy scaling model describes most simulation data within their uncertainty, despite the strong non-ideality that leads to extrema at about $x_{2}\approx 0.4\,{\rm mol\,mol^{-1}}$ (a-d in Fig. 4) or even to the occurrence of a liquid-liquid equilibrium gap (i-l in Fig. 4) – the phase behavior of these model systems was comprehensively studied in Ref. [43]. The simulation data for Ð$\textit{\DH}_{12}$ scatter more and have larger error bars than those for $D_{1}$ and $D_{2}$. The entropy scaling framework predicts the Ð$\textit{\DH}_{12}$ simulation data very well in the entire composition range. In particular, the strongly non-ideal behavior is captured accurately by the entropy scaling framework. We have compared the entropy scaling framework presented in this work to results from the established Vignes and Darken model (see Suppl. Note 8). The entropy scaling model outperforms the empirical models. Also the temperature and pressure dependency of the diffusion coefficients are very well described by the model (cf. Fig. 4 and Suppl. Note 7, respectively). For the predictions of the Fickian diffusion coefficient, the trends are correctly predicted by the entropy scaling predictions, but deviations are observed, cf. Fig. 4d,h,l. The deviations between simulation results and entropy scaling are larger for the Fickian diffusion coefficient, which can be attributed to the inclusion of the thermodynamic factor adding additional complexity to the model predictions (and the MD sampling [33]). In contrast to the other three diffusion coefficients ($D_{1}$, $D_{2}$, Ð$\textit{\DH}_{12}$), the Fickian diffusion coefficient $D_{12}$ exhibits the opposite extremum behavior, which is a result of the underlying thermodynamic factor behavior. Due to the coupling of entropy scaling and EOS, the model is inherently consistent and can be applied over a wide range of states. As an example, Fig. 4 (i-j) includes a system with weak cross-interactions resulting in a liquid-liquid equilibrium at the considered conditions. The entropy scaling model proposed in this work is able to predict not only the different diffusion coefficients in the bulk phases, but also describes the diffusion coefficients of the coexisting phases of the liquid-liquid equilibrium. Also, at the upper critical solution point, the thermodynamic factor becomes zero and so does the Fickian diffusion coefficient ($x_{2}\approx 0.38\,{\rm mol\,mol^{-1}}$). This is as expected for the Fickian diffusion coefficient at a critical point [40] and correctly captured by the model.
@@
Figure 4: Diffusion coefficients in three binary Lennard-Jones mixture with $\sigma_{2}=\sigma_{1}$ as a function of the mole fraction $x_{2}$ at $p=0.13\,\sigma^{3}_{1}\varepsilon_{1}^{-1}$ (a-d) and $p=0.26\,\sigma^{3}_{1}\varepsilon_{1}^{-1}$ (e-l). Energy parameters $\varepsilon_{1},\varepsilon_{2}$ and mixing parameter $\xi_{12}$ are given for each column. a,e,i) Self-diffusion coefficient of component 1 $D_{1}$; b,f,j) Self-diffusion coefficient of component 2 $D_{2}$; c,g,k) Maxwell-Stefan diffusion coefficient Ð$\textit{\DH}_{12}$; d,h,l) Fickian diffusion coefficient $D_{12}$. Lines are predictions from the entropy scaling framework. Symbols are simulation results from Ref. [42]. The entropy scaling framework was used in combination with the Kolafa-Nezbeda EOS[39]. The colors indicate the temperature $T\in\{0.79,0.855,0.92,0.985,1.05\}\,\rm k_{B}{}\varepsilon_{1}^{-1}$ (from blue to yellow). The black lines indicate the liquid-liquid equilibrium and the star the critical point (only i-j). The error bars represent the simulation uncertainty given in Ref. [42]. Source data are provided as a Source Data file.
@@
Fig. 5 shows the results for the self-diffusion coefficients in three real substance systems $n$-hexane + $n$-dodecane, acetone + chloroform, and nitrobenzene + $n$-hexane. Fig. 6 shows entropy scaling predictions for the Fickian diffusion coefficient in four real substance systems: toluene + $n$-hexane, acetone + chloroform, and toluene + acetonitrile. For all systems, only very limited experimental data are available (this is common for practically all real mixtures), which is used for the validation of the predictions. The thermodynamic properties of the systems were modeled by the PC-SAFT EOS [38] and the component-specific models for the pure substances were taken from the literature (see Suppl. Note 3 for details). The Berthelot combination rule parameter $\xi_{12}$ for the systems acetone + chloroform and nitrobenzene + $n$-hexane were adjusted to match the respective vapor-liquid equilibria qualitatively.
@@
Figure 5: Self-diffusion coefficients $D_{i}$ of real substance mixtures predicted by the entropy scaling model as a function of the mole fraction $x_{2}$ at $p=0.1\,{\rm MPa}$. Symbols are experimental data from the literature and lines are model predictions. a) Mixture $n$-hexane (1) + $n$-dodecane (2); experimental data from Ref. [44]. b) Mixture acetone (1) + chloroform (2); experimental data from Ref. [45]. c) Mixture nitrobenzene (1) + $n$-hexane (2); experimental data from Ref. [46]. Source data are provided as a Source Data file.